average age of 6 children is 7 years .the average age increased by 5 years when the age of father is also included.find the age of father
Answer by stanbon(75887)

average age of 6 children is 7 years .the average age increased by 5 years when the age of father is also included.find the age of father
------
sum of ages of the 6 children:: 6*7 = 42
----
Equation:
Let father's age be "x"::
(42+x)/7 = 7+5
42+x = 7*12
x = 42 yrs (father's age)
--------------------------------
Cheers,
Stan H.
--------------------